# CamReasoner: Reinforcing Camera Movement Understanding via Structured Spatial Reasoning

### v2

@@ 4.1 Setup
The RL phase is implemented using the verl framework with a learning rate of $2\times 10^{-6}$ and bf16 mixed-precision. We employ a rollout batch size of 128 and a global batch size of 32. To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution. The total reward is defined as $R=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$. Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
@@ 4.2 Quantitative Results
Table 1: Binary classification performance on camera movement understanding. CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs.
{NiceTabular}
lccccccccccccccc|$c$ Model Translation (Dolly/Pedestal/Truck) Zooming Rotation (Pan/Tilt/Roll) Static Avg
@@
Random Chance 29.3 9.7 6.7 8.6 15.8 11.5 11.1 10.2 15.0 15.4 12.7 7.7 8.9 10.2 9.7 12.2
SfM/SLAM
COLMAP [schonberger2016structure] 36.2 13.1 11.9 19.7 34.1 30.0 13.9 14.2 43.9 46.4 28.3 19.1 42.1 48.7 7.5 27.3
VGGSFM [wang2024vggsfm] 56.6 28.9 28.7 38.2 48.9 35.3 21.7 17.3 60.9 58.7 46.6 43.3 61.4 55.5 16.7 41.3
DUSt3R [wang2024dust3r] 58.9 24.0 30.7 18.0 38.3 26.9 18.2 24.6 59.4 63.8 32.9 27.3 61.0 57.9 13.1 37.0
MASt3R [duisterhof2024mast3r] 47.5 21.1 23.5 40.2 38.7 38.1 42.2 46.6 66.6 58.0 63.2 40.3 50.4 53.5 15.7 43.1
CUT3R [wang2025continuous3dperceptionmodel] 68.9 50.4 24.7 34.2 37.0 27.6 15.9 21.3 59.1 65.0 65.0 47.5 60.7 66.2 15.1 42.7
MegaSAM [li2024megasam] 73.8 43.9 24.2 29.1 45.3 44.2 11.1 10.2 79.5 82.2 73.8 65.3 71.5 75.8 22.0 50.1
VQAScore
LLaVA-OneVision-7B [li2024llavaonevision] 48.2 13.9 13.0 17.4 24.4 20.8 11.0 14.8 34.5 34.6 17.4 32.3 19.9 21.4 19.4 22.9
LLaVA-Video-7B [zhang2025llavavideo] 56.3 15.7 17.0 19.9 27.9 24.3 16.7 17.4 34.6 37.9 27.7 38.3 16.6 22.4 22.8 26.4
InternVideo2-Chat-8B [wang2024internvideo2] 72.0 19.1 19.9 18.1 18.4 24.1 12.6 10.7 23.3 23.4 17.7 23.5 20.2 16.9 20.8 22.7
Tarsier-Recap-7B [yuan2025tarsier2] 61.5 15.6 26.5 24.4 29.7 22.1 14.8 15.5 23.5 28.1 25.3 22.2 15.7 19.3 31.6 21.6
InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 50.5 10.9 11.7 10.7 15.0 10.9 12.2 17.0 14.7 14.3 15.1 18.0 12.1 18.6 22.5 17.0
InternVL2.5-8B [chen2025expandingperformance] 69.9 13.3 28.9 26.7 24.1 23.9 19.2 33.1 38.5 31.8 38.7 38.0 11.8 26.1 24.1 30.4
InternVL2.5-26B [chen2025expandingperformance] 65.5 12.2 21.7 24.3 28.0 20.0 22.5 32.5 43.8 39.4 46.2 44.9 14.7 18.7 25.9 30.7
GPT-4o [openai2024gpt4ocard] 68.3 30.1 21.7 39.3 39.1 22.6 43.0 40.5 46.0 43.4 44.9 36.6 24.7 29.6 33.0 37.5
InternVL3-8B [zhu2025internvl3e] 63.0 16.0 19.4 29.9 31.4 28.1 30.4 28.9 42.8 50.8 43.3 37.6 21.9 23.0 20.7 32.4
InternVL3-78B [zhu2025internvl3e] 74.2 18.7 20.2 33.5 34.8 30.3 27.2 34.4 48.6 55.1 49.2 41.5 28.4 25.8 23.3 37.9
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 57.7 15.3 19.3 31.4 35.5 28.4 30.7 44.7 64.6 68.7 56.1 35.1 19.4 24.9 20.4 36.9
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 59.3 16.9 20.8 33.1 37.1 30.1 32.3 46.4 66.2 70.2 57.7 36.7 20.8 26.5 21.8 38.4
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 59.7 17.3 21.2 33.5 37.5 30.4 32.7 46.8 66.6 70.7 58.1 37.1 21.2 26.9 22.2 38.8
SOTA Baseline
Cam-Motion-7B [lin2025towards] 86.4 40.5 29.2 50.2 71.3 52.5 57.2 79.6 83.2 87.1 80.1 71.0 33.9 43.1 77.3 63.0
Cam-Motion-32B [lin2025towards] 88.2 42.1 30.8 51.9 73.1 54.1 58.8 81.2 84.7 88.8 81.8 72.7 35.6 44.8 79.2 64.7
Cam-Motion-72B [lin2025towards] 89.4 43.4 32.0 53.1 74.2 55.2 60.0 82.4 86.0 90.1 83.1 73.9 36.8 46.0 80.4 65.9
@@
As presented in Tab. 4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models. Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o and Qwen2.5-VL-72B but also yields a substantial improvement over the task-specific SFT baseline. This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks. Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds. Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the performance of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines. By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting. Consequently, these results validate that our framework successfully bridges the gap between raw visual perception and high-level camera movement reasoning, providing a robust foundation for automated cinematography analysis.
Table 2: VQA evaluation on the CameraBench. We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions with questions on camera movement type in given video clips. Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
{NiceTabular}
@@
Random Chance 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
mPLUG-Owl3-7B 53.1 66.4 62.9 49.8 50.4 55.4 54.5 47.1 64.9 57.2
LLaVA-Video-7B [zhang2025llavavideo] 54.8 67.7 58.6 53.4 51.1 56.3 61.3 52.6 69.6 60.2
LLaVA-OneVision-7B [li2024llavaonevision] 55.7 65.3 70.6 54.4 56.8 62.4 62.2 44.4 53.6 58.5
InternVideo2-Chat-8B [wang2024internvideo2] 53.7 65.9 53.0 51.5 51.0 53.5 49.7 52.2 51.8 52.6
Tarsier-Recap-7B [yuan2025tarsier2] 53.1 64.3 51.7 51.0 50.2 52.8 49.0 51.4 51.0 51.8
InternVL2.5-8B [chen2025expandingperformance] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
InternVL2.5-26B [chen2025expandingperformance] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
InternVL3-8B [zhu2025internvl3e] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
InternVL3-78B [zhu2025internvl3e] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
GPT-4o [openai2024gpt4ocard] 57.2 53.9 62.7 59.5 54.6 65.6 53.0 43.1 63.4 60.4
Gemini-2-Flash [comanici2025gemini25] 54.9 47.9 58.0 45.6 42.1 47.6 47.6 40.2 65.3 53.1
Gemini-2.5-Pro [comanici2025gemini25] 59.6 52.5 61.5 50.1 46.8 53.6 50.9 43.9 66.1 56.1
Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport] 59.3 58.3 60.6 54.0 51.7 55.7 57.6 50.0 66.7 59.2
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.8 60.6 62.9 56.9 54.0 58.4 64.2 53.6 64.4 60.3
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 57.7 62.1 71.0 57.8 53.2 62.4 59.2 50.4 68.6 60.3
@@
These findings are further corroborated by the VQA evaluation results presented in Tab. 4.2. CamReasoner-7B (Ours) maintains its dominance across a diverse array of evaluative dimensions, achieving a superior overall average accuracy of 74.5%. Our model establishes new performance ceilings in critical categories such as Motion & Steadiness (78.0%), Scene Dynamics (77.0%), and Shot Tracking (71.3%). Notably, the model demonstrates exceptional proficiency in identifying complex scenarios, reaching 82.6% in Complex Description and 60.7% in Confusable Motion, significantly outperforming previous state-of-the-art results. Furthermore, it shows high precision in logical reasoning tasks, as evidenced by its performance in Has Motion (71.8%) and Only Motion (60.6%) categories. This consistent outperformance across both binary classification and multidimensional VQA tasks reinforces the model’s robust capacity for high-fidelity camera movement understanding.
@@ 4.3 Qualitative Results
During the observation phase, the model performs a detailed temporal scan to capture subtle shifts in subject-background relations. This granular perception allows it to decouple ego-motion from environmental noise, explicitly describing transitions between visual landmarks. In the subsequent thinking stage, CamReasoner engages in deductive verification grounded in cinematographic principles. By cross-referencing perceived displacements with the absence of confounding rotations, the model demonstrates that it has internalized robust spatial-temporal paradigms rather than relying on superficial statistical correlations.
@@ 4.4 Training Curves
The training dynamics of the SFT phase exhibit robust convergence and stability, as evidenced by the synchronized progression of loss, gradient norm, and learning rate in Figure 4. Specifically, the training loss undergoes a rapid initial decline and follows a distinct step-wise reduction at each epoch boundary, eventually stabilizing at approximately 0.05, which indicates that the model has effectively captured the underlying distribution of camera movement reasoning tasks. Throughout this process, the gradient norm remains within a well-behaved range between 2.0 and 4.0, ensuring a stable optimization landscape under the peak learning rate of $1.0\times 10^{-5}$ and its subsequent cosine decay. Notably, we select the checkpoint from the conclusion of the first epoch as the initialization for the subsequent RL stage; at this juncture, the model has successfully acquired the fundamental paradigms of spatial reasoning and the <observation>-<think>-<answer>output format, yet retains sufficient policy entropy by avoiding the potential overfitting characteristic of later training stages.
The RL phase, initialized from the first-epoch SFT checkpoint, further refines the model’s policy to achieve rigorous logical alignment. As illustrated in Figure 4, the training dynamics demonstrate a consistent upward trend across all reward metrics. Specifically, the reward/format metric experiences a rapid surge within the first 20 steps, quickly stabilizing near a perfect score of 1.0. This indicates that the model swiftly masters the <observation>-<think>-<answer>structural constraint, even under the exploration pressure of the RL environment. Concurrent with format stabilization, the reward/accuracy exhibits a steady and robust climb from approximately 0.68 to over 0.85, proving that the GRPO-based optimization effectively sharpens the model’s camera-centric reasoning boundaries. The synchronized growth of the reward/overall metric, which eventually plateaus at a high equilibrium, confirms the successful balance between linguistic consistency and spatial-temporal accuracy. As shown in the qualitative results, this optimized policy enables CamReasoner-7B to execute meticulous deductive chains—such as explicitly verifying the absence of confounding motions in complex Truck or Pedestal shots—thereby bridging the gap between raw perception and logical cinematic analysis.

### v3

@@ 4.1 Experimental Setup
We evaluate all models under a unified protocol: video input is standardized to a maximum of 64 frames and 131,072 pixels ($128\times 32\times 32$) sampled at 1 FPS. The SFT phase uses LLaMA-Factory under DeepSpeed ZeRO-3, freezing the vision tower and multi-modal projector while fully fine-tuning the language model at a learning rate of $1\times 10^{-5}$ with a cosine schedule for 3 epochs. The RL phase is implemented with verl at a learning rate of $2\times 10^{-6}$ for 2 epochs, sampling n=8n{=}8 rollouts per prompt with a low-variance KL penalty of $\beta{=}1\times 10^{-2}$ and online filtering that prunes the top/bottom 1% of reward samples. We set the $\lambda$ in Eq. 3 to 0.1 to balance different rewards. The policy is initialized from the second SFT epoch to balance reasoning capacity with exploration entropy. Full configurations and computational cost are in Appendix C.
@@ 4.2 Quantitative Results
Table 1: Binary classification performance on camera movement understanding. Built upon Qwen2.5-VL-7B, CamReasoner-7B demonstrates significant improvements over its backbone and outperforms both proprietary and open-source baselines across most categories. Best results are in bold and second best are underlined.
Model     Translation (Dolly/Pedestal/Truck)     Zooming     Rotation (Pan/Tilt/Roll)     Static     Avg
@@
GPT-4o [openai2024gpt4ocard]     63.6 61.6 67.2 53.5 69.6 59.4     76.0 57.1     64.5 76.2 63.1 77.0 60.0 74.5     74.7     66.5
@@
LLaVA-Video-7B [zhang2025llavavideo]     35.5 46.1 31.4 53.5 48.9 41.2     48.4 22.6     15.2 21.9 20.3 23.7 44.7 59.6     32.9     37.7
InternVL2.5-8B [chen2025expandingperformance]     44.3 55.2 61.9 66.0 46.4 66.0     61.8 65.3     63.9 55.3 68.0 64.5 63.6 50.9     52.0     59.0
Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport]     64.7 61.6 84.7 81.9 71.3 79.9     85.1 72.3     77.9 79.5 84.9 89.7 71.7 75.5     17.8     69.3
InternVL2.5-26B [chen2025expandingperformance]     62.1 50.8 75.3 80.5 74.0 72.8     58.0 61.5     72.0 71.4 82.0 78.8 66.8 65.2     71.8     69.5
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport]     62.7 56.8 82.8 82.8 75.8 79.7     58.7 68.6     72.9 80.1 83.8 86.0 67.9 73.8     74.2     73.8
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport]     73.6 76.1 80.2 80.2 72.5 65.2     67.6 73.9     78.5 78.9 71.6 66.7 89.4 78.3     78.0     75.4
[…]
Cam-Motion-7B [lin2025towards]     39.7 48.6 32.6 64.2 50.5 54.5     51.3 14.2     17.6 18.0 12.0 20.8 47.2 75.1     35.0     38.8
Cam-Motion-72B [lin2025towards]     69.4 79.2 80.0 79.5 77.6 77.5     81.8 80.3     80.5 80.7 84.7 87.5 84.5 82.3     82.8     80.6
[…]
Binary Classification.
As presented in Tab. 4.2, CamReasoner-7B achieves an overall accuracy of 78.4%, surpassing GPT-4o by 11.9 percentage points and Qwen2.5-VL-32B by 3.0 points using a 7B backbone, and approaching Cam-Motion-72B within 2.2 points despite a 10× parameter difference. Compared to its backbone Qwen2.5-VL-7B, CamReasoner-7B yields a consistent average gain of 4.6 points, with the most pronounced improvements in Rotation and Zooming categories: Zoom In increases from 58.7% to 74.2%, Pan Right from 72.9% to 85.7%, and Move Right from 75.8% to 84.6%. Static recognition similarly improves from 74.2% to 80.1%, indicating enhanced capacity to classify stationary cameras in the presence of dynamic scene content. We further observe that the backbone exhibits a 4.0-point gap between Dolly In and Zoom In, suggesting limited ability to discriminate between the two visually similar approaching motions. Following O-T-A training, both categories improve to 68.7% and 74.2% respectively, with the gap reversed, which we attribute to the explicit spatial reasoning process encouraging finer-grained analysis of motion characteristics. A marginal regression is observed in Roll CCW, which decreases from 73.8% to 70.2%, suggesting that the structured reasoning process may introduce slight uncertainty in rotation directions where the backbone already performs relatively well.
Table 2: VQA evaluation on CameraBench across nine fine-grained dimensions of camera movement understanding. Built upon Qwen2.5-VL-7B, CamReasoner-7B consistently improves over its backbone across all categories and outperforms both proprietary and open-source baselines. Best results are in bold and second best are underlined.
Model     Motion & Scene Motion Motion Confusable Has Shot Only Complex     Avg
@@
GPT-4o [openai2024gpt4ocard]     61.3 57.2 66.8 62.1 57.3 68.4 58.5 47.6 67.8     64.5
[…]
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport]     46.5 33.6 52.3 52.8 38.0 42.4 41.9 44.4 46.8     45.3
LLaVA-Video-7B [zhang2025llavavideo]     42.8 44.3 46.8 48.4 26.7 54.9 52.4 34.0 59.3     52.3
InternVL2.5-8B [chen2025expandingperformance]     44.1 55.2 49.1 51.5 27.3 56.0 53.9 36.3 60.6     54.1
Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport]     51.0 67.2 76.4 56.4 40.7 58.3 58.1 45.3 62.8     58.6
InternVL2.5-26B [chen2025expandingperformance]     56.4 39.4 54.3 62.9 36.0 56.2 61.1 60.6 63.3     58.9
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport]     59.3 48.0 65.2 64.9 48.7 57.2 63.5 64.2 63.6     60.9
[…]
Cam-Motion-7B [lin2025towards]     45.3 54.0 46.3 23.2 21.0 46.9 50.7 36.2 37.6     40.5
Cam-Motion-72B [lin2025towards]     53.1 60.9 48.3 52.4 62.7 50.1 49.5 58.1 50.7     51.6
[…]
VQA Evaluation.
To further assess the model’s ability to reason over competing motion descriptions, we evaluate on the VQA task under our reformulated multiple-choice protocol. As shown in Tab. 4.2, CamReasoner-7B achieves 74.5% overall, improving over its backbone Qwen2.5-VL-7B by 13.6 percentage points and exceeding GPT-4o by 10.0 points. The Cam-Motion models score between 40.5% and 51.6%, notably below their binary classification results. As discussed in Appendix B, the original CameraBench evaluates VQA via pairwise confidence-score comparison under a binary yes/no protocol, and the Cam-Motion models are optimized for this setting. Our reformulated protocol instead requires the model to select among competing motion descriptions, assessing explicit reasoning rather than relative score calibration. This shift inevitably disadvantages models trained exclusively under the original protocol, so we anchor our VQA analysis on comparisons with models that natively support multiple-choice output. Under this protocol, CamReasoner-7B demonstrates substantial improvements across most dimensions. The largest gains over the backbone appear on Scene Dynamics, rising from 48.0% to 77.0%, Complex Description, from 63.6% to 82.6%, and Motion & Steadiness, from 59.3% to 78.0%. The improvement on Complex Description is particularly noteworthy, as the structured decomposition in the observation stage and synthesis in the thinking stage enable joint reasoning over multiple co-occurring motion attributes. The accuracy on Confusable Motion at 60.7% and Only Motion at 60.6%, while among the highest across all compared models, remains relatively modest, reflecting the inherent difficulty of distinguishing visually similar motion pairs under the multiple-choice setting.
Out-of-domain Generalization and Reasoning Reliability.
To assess whether the acquired reasoning transfers beyond the training distribution, we evaluate on ShotBench [liu2025shotbench] and RefineShot [wu2025refineshot], neither of which is included in our training data. As shown in Tab. 4.2, CamReasoner-7B achieves 39.7%, improving over Qwen2.5-VL-7B at 30.2% by 9.5 percentage points. While the absolute accuracy reflects the challenging nature of ShotBench, the consistent improvement over the backbone suggests that the structured reasoning acquired through O-T-A training is not limited to the CameraBench distribution. We note that InternLM-XC2.5-7B at 38.8% and InternVL3-14B at 38.2% achieve comparable accuracy without in-domain training, indicating that further improving out-of-domain generalization remains an open challenge. Regarding reasoning reliability, CamReasoner-7B scores 96.1%, slightly below the backbone’s 98.9%. This is expected, as the O-T-A paradigm produces substantially longer outputs, approximately 200 tokens versus fewer than 20 for direct-answer baselines, which increases the chance of minor inconsistencies across semantically equivalent queries. The reliability nevertheless remains comparable to InternVL3-14B at 97.1%, confirming that the longer reasoning chains do not compromise overall logical coherence.
Table 3: Performance and reliability on ShotBench and RefineShot. CamReasoner achieves a +9.5% accuracy gain over its Qwen2.5VL-7B backbone while maintaining high reasoning reliability (96.1%).
Model     Size     Accuracy Reliability Score
[…]
Iblip-vicuna-7B [dai2023instructblip]     7B     25.0 -
Iblip-vicuna-13B [dai2023instructblip]     13B     22.0 -
VILA1.5-3B [lin2024vila]     3B     21.5 -
VILA1.5-8B [lin2024vila]     8B     36.9 -
VILA1.5-13B [lin2024vila]     13B     31.3 -
LLaVA-NeXT-7B [li2024llava]     7B     31.3 -
LLaVA-VID-7B [li2024llama]     7B     35.3 -
Ovis2-8B [lu2024ovis]     8B     35.3 -
InternLM-XC2.5-7B [dong2024internlm]     7B     38.8 -
InternVL2.5-8B [chen2025expandingperformance]     8B     33.6 96.3
InternVL3-2B [zhu2025internvl3e]     2B     38.1 95.4
InternVL3-14B [zhu2025internvl3e]     14B     38.2 97.1
[…]
Qwen2.5VL-3B [bai2025qwen25vltechnicalreport]     3B     31.9 98.5
Qwen2.5VL-7B [bai2025qwen25vltechnicalreport]     7B     30.2 98.9
[…]
@@ 4.3 Qualitative Results
During the observation phase, the model performs a detailed temporal scan to capture subtle shifts in subject-background relations. This granular perception allows it to decouple ego-motion from environmental noise, explicitly describing transitions between visual landmarks. In the subsequent thinking stage, CamReasoner engages in deductive verification grounded in cinematographic principles. By cross-referencing perceived displacements with the absence of confounding rotations, the model demonstrates that it has learned to produce structured reasoning chains that reference relevant spatial-temporal cues rather than relying on superficial statistical correlations.
@@ 4.4 Training Curves
The training dynamics of the SFT phase exhibit robust convergence and stability, as evidenced by the synchronized progression of loss and gradient norm in Figure 4. Specifically, the training loss undergoes a rapid initial decline and follows a distinct step-wise reduction at each epoch boundary, eventually stabilizing at approximately 0.05, which indicates that the model has effectively captured the underlying distribution of camera movement reasoning tasks. Throughout this process, the gradient norm remains within a well-behaved range between 2.0 and 4.0, ensuring a stable optimization landscape under the peak learning rate of $1.0\times 10^{-5}$ and its subsequent cosine decay. Notably, we select the checkpoint from the conclusion of the second epoch as the initialization for the subsequent RL stage; at this juncture, the model has successfully acquired the fundamental paradigms of spatial reasoning and the <observation>-<think>-<answer>output format, yet retains sufficient policy entropy by avoiding the potential overfitting characteristic of later training stages.
Table 4: Ablation studies on CamReasoner. We evaluate the contribution of each training stage and key RL hyperparameters.
(a) Training stages and RL initialization timing. Bin. and VQA denote binary classification and visual question answering accuracy (%), respectively.
Model SFT RL Bin. Avg. VQA Avg.
Base (Qwen2.5-VL-7B) 73.8 60.9
SFT at different epochs:
Epoch. 1 74.5 65.8
Epoch. 2 75.3 69.5
Epoch. 3 75.8 69.2
RL from different SFT checkpoints:
Epoch. 1 $\rightarrow$ RL 76.2 71.2
Epoch. 2 $\rightarrow$ RL (Ours) 78.4 74.5
Epoch. 3 $\rightarrow$ RL 76.5 70.7
(b) RL hyperparameters: reward weight $\lambda$, GRPO group size $G$, and advantage normalization (standard (Std.) vs. EMA-GRPO with momentum $\alpha$).
Reward weight $\lambda$ 0.0 0.1 0.3 0.5
Bin. avg. 74.6 78.4 75.9 71.4
VQA avg. 70.2 74.5 72.1 68.3
GRPO group size $G$ 2 4 8 16
Bin. avg. 73.8 75.8 78.4 78.1
VQA avg. 70.6 72.3 74.5 74.2
Adv. normalization Std. $\alpha$=0.9 $\bm{\alpha}$=0.99 $\alpha$=0.999
Bin. avg. 75.1 76.8 78.4 77.6
VQA avg. 71.5 73.0 74.5 73.8
The RL phase, initialized from the second-epoch SFT checkpoint, further refines the model’s policy to achieve rigorous logical alignment. As illustrated in Figure 4, the training dynamics demonstrate a consistent upward trend across all reward metrics. Specifically, the reward/format metric experiences a rapid surge within the first 20 steps, quickly stabilizing near a perfect score of 1.0. This indicates that the model swiftly masters the <observation>-<think>-<answer>structural constraint, even under the exploration pressure of the RL environment. Concurrent with format stabilization, the reward/accuracy exhibits a steady and robust climb from approximately 0.68 to over 0.85, proving that the GRPO-based optimization effectively sharpens the model’s camera-centric reasoning boundaries. The synchronized growth of the reward/overall metric, which eventually plateaus at a high equilibrium, confirms the successful balance between linguistic consistency and spatial-temporal accuracy. As shown in the qualitative results, this optimized policy enables CamReasoner-7B to execute meticulous deductive chains—such as explicitly verifying the absence of confounding motions in complex Truck or Pedestal shots—thereby bridging the gap between raw perception and logical cinematic analysis.